OpenStudy (anonymous):

Hello! may I have help with how to begin this? Use Newton's method to find all roots of the equation correct to six decimal places. x^4 = 9 + x

OpenStudy (dumbcow):

set equal to 0 find derivative set value for initial guess

OpenStudy (anonymous):

@gamalavolti can you start?

OpenStudy (anonymous):

x^4 = 9 + x x^4 - x - 9 = 0

OpenStudy (anonymous):

@Gmalavolti